D)Suppose you deposit $2,000 for 5 years at a rate of 8%.
a) Calculate the return (A) if the bank compounds annually (n=1). Round your answer to the nearest cent.
b) Calculate the return (A) if the bank compounds quarterly (n=4).
c) Does compounding annually or quarterly yield more interest?
d) If a bank compounds continuously, then the formula used is A=Pe^rt where e is a constant and equals approximately 2.7183. Calculate A with continuous compounding. Round your answer to the nearest cent.
e) A commonly asked question is, "How long will it take to double my money?" At 8% interest rate and continuous compounding, what is the answer?

Answer by rfer(16322)

A)
A=P(1+r)^nt
A=2000(1+.08)^5
A=2000(1.46933)
A=$2938.65
B)
A=2000(1+.08/4)^4*5
A=2000(1.02)^20
A=2000(1.485947)
A=$2971.89
C) quarterly
D)
A=Pe^rt
A=2000(2.7183)^.08*5
A=2000(1.491828687)
A=$2983.67
E)
72/8%=9yrs